OpenStudy (anonymous):

Probability: In how many different ways can a 10 question true-false test be answered if every question must be answered?

OpenStudy (nurali):

each question 2 choices => 2*2* 2*2*2*2* 2*2*2*2 = 2^10 = 1024

OpenStudy (anonymous):

all right, thanks ! :)